OpenStudy (anonymous):

find dy/dx given y = definite integral 1/(1+t^2). upper cosx. lower 0. show steps/explain.

OpenStudy (anonymous):

\[\Large \frac{dy}{dx} \int_0^{\cos x} \frac{1}{1+t^2}dt \] This one?

OpenStudy (anonymous):

Seems like a problem for the fundamental theorem of Calculus.

OpenStudy (anonymous):

well it says y = that integral. find dy/dx.

OpenStudy (anonymous):

\[ \Large \int_0^x =F(x)-F(0)\] If you derive the above you get again \[ \Large \frac{dy}{dx} \int_0^x=f(x) \]

OpenStudy (anonymous):

So I would say, just substitute cosx into t^2

OpenStudy (anonymous):

ya, but what is the numerator

OpenStudy (anonymous):

\[\Large \frac{dy}{dx} \int_0^{\cos x}=f(\cos x) = \frac{1}{1+\cos^2x}\]

OpenStudy (anonymous):

that's the correct denom...

OpenStudy (anonymous):

what's the solution then?

OpenStudy (turingtest):

gotta have a -sinx in the top, eh?

OpenStudy (anonymous):

im not sure... i think -sinx

OpenStudy (anonymous):

hmm so I forgot about the dt.

OpenStudy (anonymous):

but I wonder why that would not fall out of the solution anyway, because it's differentiated.

OpenStudy (anonymous):

ah, I see. they want it like this? \[\Large dy \int_0^{\cos x}=f(\cos x) dx \] ?

OpenStudy (turingtest):

\[\Large \frac{d}{dx}\left(\int_0^{g(x)}f(t)dt\right)=\]\[\frac d{dx}\left(F(g(x))-\cancel{F(0)}^{\large0} \right)= F'(g(x))g'(x)=f(x)g'(x)\]chain rule

OpenStudy (anonymous):

i dunno, but im satisfied with -sinx / 1+cos^2x

OpenStudy (turingtest):

I repeat, chain rule the derivative is with respect to x, and our upper bound is a function of x

OpenStudy (anonymous):

oh that's where I went wrong.

OpenStudy (anonymous):

ty both!

OpenStudy (turingtest):

in general the fundamental theorem of calculus would be\[\Large \frac{d}{dx}\left(\int_{g(x)}^{h(x)}f(t)dt\right)\]\[\large =\frac d{dx}\left(F(h(x))-F(g(x))\right)= F'(h(x))h'(x)-F'(g(x))g'(x)\]where\[F'(x)=f(x)\]

OpenStudy (anonymous):

\[ \Large \int_0^{\cos x} =F(\cos(x))-F(0) \] Would have been best if I had made this 'extra' step to make sure about the chain rule \[ \Large \frac{dy}{dx} \Large \int_0^{\cos x}=-f(\cos(x))\sin(x) \]

OpenStudy (turingtest):

I wouldn't write the y in the dy/dx, since y is the integral itself and is already present in the equation

OpenStudy (anonymous):

Yes you're right, standard Leibniz Notation? \[ \Large \frac{d}{dx} \Large \int_{0}^{\cos x}f(t)dt=-f(\cos(x))\sin(x) \]

OpenStudy (turingtest):

now that is exact, yeah

OpenStudy (anonymous):

For the above case of course, the boundaries are depending on y.

OpenStudy (anonymous):

okay thanks a lot!

OpenStudy (turingtest):

welcome!